OpenStudy (anonymous):

To approach the runway, a pilot of a small plane must begin a 10 descent starting from a height of 1983 feet above the ground. To the nearest tenth of a mile, how many miles from the runway is the airplane at the start of this approach?

OpenStudy (anonymous):



OpenStudy (anonymous):

@OpenSessame

OpenStudy (anonymous):

sin10=1983/d

OpenStudy (anonymous):

BTW 5,280 feet in one mile

OpenStudy (anonymous):

my answer choices are A. 2.2 mi B. 0.4 mi C. 11,419.6 mi D. 2.1 mi

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

try solving it! with what i gave you

OpenStudy (anonymous):

I don't understand how to solve it like that

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

D=1983(sin10)

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

SORRY

OpenStudy (anonymous):

do you have a calculator?

OpenStudy (anonymous):

if you do type in sin10

OpenStudy (anonymous):

then equal.

OpenStudy (anonymous):

then multiply that by 1983

OpenStudy (anonymous):

I got 344.3

OpenStudy (anonymous):

now thats in feet you have to put it in miles

OpenStudy (anonymous):

so if I convert the 344.3 from feet to miles would I do 344.3* 5,280

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

when I divided the two I got 15.3

OpenStudy (anonymous):

one sec...

OpenStudy (anonymous):

did you so the first part right?

OpenStudy (anonymous):

okay you lost me

OpenStudy (anonymous):

sin(10 degrees) = 1983/x x = 1983/sin(10 degrees) x = ? ft.

OpenStudy (anonymous):

its D...

OpenStudy (anonymous):

that was the wrong answer @OpenSessame

OpenStudy (anonymous):

really? cause i cant figure this out.

OpenStudy (anonymous):

yeah I got it graded and it was counted wrong

OpenStudy (anonymous):

can you take it again?

OpenStudy (anonymous):

no but its okay thanks for all the help. if I have any questions in the future would it be okay to tag you in it to see if I can get your help?

OpenStudy (anonymous):

yeap:) are you doing ashworth?

OpenStudy (anonymous):

Well wouldnt you guess the answer is 2.2....*facepalm*

OpenStudy (anonymous):

oh well it's okay and yeah I'm doing ashworth

OpenStudy (anonymous):

Me too!